40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4862     Entry time: Thu Jun 23 02:12:12 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: work schedule. 

 June 22-June 24:

1.Coupling light into fibre at the ETMY.

2.Routing of the fibre to the PSL table.

June 27-June 30:

1.PSL optical table layout sketching.

2.Combining the PSL beam with fibre output onto a BS and then superpose them on a New Focus 1611 PD.

July 5-July 8:

1.Conversion of the PD output to voltage using MFD(Mixer Frequency Discriminator).

2. Report writing.

July 7: 5:00 pm: 1st Report Due.

July 11-July 22:

1.Locking Y-arm to PSL.

2.Setting up the feedback loop using the MFD output as the error signal and acting on the AUX laser frequency.

July 25-Aug 5:

1.Y-Arm cavity characterisation.

Measurement of the transmission of IR and green light through the cavity.

2.Analysis.

To obtain FSR, Finesse,Loss of the Cavity, Visibility, Transverse Modes(g-factor, astigmatism), Reflectivity, Q-factor.

3.Report and abstract writing.

Aug 1: 5:00 pm: 2nd Report and absract due.

Aug 8-11:

Preparation for talk and seminar.

ELOG V3.1.3-